OpenStudy (anonymous):

Solve for y. 4x – 1 = 6 + 3y + 3x x + 4 + y = 11 + 4y

OpenStudy (arabpride):

1) 4x – 1 = 6 + 3y + 3x combine like terms: 4x – 1 = 6 + 3y + 3x -3x -3x x - 1 = 6 + 3y +1 +1 x = 3y + 7 (this is simplified) 2) x + 4 + y = 11 + 4y combine like terms: x + 4 + y = 11 + 4y -y -y x + 4 = 11 + 3y -4 -4 x = 3y + 7 (simplified)

OpenStudy (anonymous):

Thank you!